# shapes.geometric, shapes.multipart, arrows, arrows.meta, positioning, calc, chains, backgrounds Towards Explainable Evaluation Metrics for Machine Translation

### v1

@@ 2.2 Explainability
Explanation type Further, we distinguish explanation types, i.e., the output format of the explanations (e.g. Wiegreffe and Marasovic 2021). One example are natural language explanations that describe the model output or its internal decision process with free text. In the case of MT metrics, this could be a sentence like “The metric returned a score of 0.6 for the translation, as it correctly captures all details of the source sentence but has severe grammatical errors at positions […]”. There are many possible explanation types (e.g. Barredo Arrieta et al. 2020; Wiegreffe and Pinter 2019), such as, feature importance explanations or counterfactuals. For our taxonomy, we select those that occur in our target papers (see §3). The definition of each explanation type is paired with the description of methods in §4.
@@
Model Access Most taxonomies of explainability techniques consider the required level of model access (e.g. Barredo Arrieta et al. 2020; Linardatos et al. 2021; Madsen et al. 2022). Here, techniques that require no access (that can work with any black box model) are called model-agnostic. Techniques that require access to model parameters are called model-specific, as only specific model architectures satisfy the requirements of the method. Due to the prevalent role of transformer architectures, many modern evaluation metrics use embeddings and attention mechanisms and allow for gradient computations. This allows to apply many techniques that are specific to these architectures to such metrics. A special type of model-specific explanations are those that are given by a model (here a metric) itself. We call these metrics “explainable by design”, as they are designed in a way that provides more details on their explanations or decision process. Examples for MT metrics in this category are unsupervised matching metrics like BERTScore Zhang et al. (2020a) and MoverScore Zhao et al. (2019), as they compute token-level similarity scores that can be used as explanations of the sentence-level score Leiter (2021); Fomicheva et al. (2021). Finally, there are also interpretable models which are easy to understand for humans such as Naive Bayes and decision trees Freitas (2014); Arya et al. (2019). We can normally obtain decision-understanding explanations of these models at the same time as predictions (for example, the corresponding path in the decision tree), while the models can also be considered explanations of themselves (for example, the trained decision tree itself). We do not use the dimension of interpretable models in our taxonomy (see §4) as none of the state-of-the-art metrics falls into this category. Model-agnostic techniques require least model access, model-specific techniques, which use the gradient, embeddings or attention weights require more access and models that are explainable by design require access to the full model to generate explanations.888Note that besides interpretable models, none of the access levels inherently leads to faithful explanations, i.e. explanations that correctly describe the internal processes of a model. Hence, this property needs to be considered separately (e.g. Jacovi and Goldberg 2020).
@@
Audiences Prior works in explainability point out that different audiences require different types of explanations (e.g. Barredo Arrieta et al. 2020; Sharma et al. 2020). For example, non-experts might have difficulties understanding explanations given in the domain language of experts. Barredo Arrieta et al. (2020) distinguish model users, affected users, developers, regulatory entities and managers. Users of MT metrics are usually domain experts in MT that want to evaluate or improve MT systems, i.e. MT developers. Another group of users might be non-experts that want to use the models to check their own translations or machine generated translations. Affected users might be the ones that receive translations that were graded by a metric or generated by a model that was supervised or tested with the metric. For example, Fernandes et al. (2022a) employ MT metrics in the decoding process of MT models. We consider users of those MT models as affected users. Developers are the designers of MT metrics. Regulatory entities and managers have an interest in enforcing guidelines. We describe the relation of the audiences towards goals of explainable MT metrics in §2.3 and associate audiences with explanation types in §4.
@@ 2.3 Goals of Explainable MT Metrics
In this section, we discuss goals that require specific focus for explainable MT evaluation. We orient ourselves on common goals proposed by Lipton (2016); Barredo Arrieta et al. (2020); Jacovi and Goldberg (2021) and identify four specific ones that match the literature we identify in §3 and use cases that might commonly occur in our field of research (see Figure 3). We follow with a short description of each of them below.
@@
Goal 1: Diagnose and improve metrics. By explaining why a metric predicted a certain score for a machine translation, where humans would assign a different score, developers might understand its weaknesses (e.g. Kaster et al. 2021; Sai et al. 2021). This can enable architectural changes or the selection of different data sets, leading to metric improvement. Likewise, explaining the whole metric can unveil whether it follows desired general properties or might otherwise be led astray by carefully crafted adversarial inputs. Goal 1 is especially important for the audience metric developers, as they need to understand the shortcomings to improve on them.
@@
Goal 2: Make metrics more expressive and accessible. A metric that assigns a score on a sentence level is difficult to understand for non-experts, e.g., non-native speakers. If a metric provides further information, such as which words it considers incorrect, it may become more accessible and thus foster trust (which may lead to more widespread adoption of the metric). Exemplary works in this direction are Fomicheva et al. (2021); Bao et al. (2023). Accessible metrics might also help with use cases like translation error correction in the downstream tasks of language learning or human translation services (e.g. Zouhar et al. 2023). In these latter two use cases, accessibility might not require faithfulness of the explanation (see §2.2). Goal 2 is important for the audience of metric users. These can be either non-experts that want to understand specific metric outputs or developers of MT systems that want to understand if a metric is a good fit for their tasks (see Figure 1).
@@
Goal 3: Support Semi-Automatic labeling. Fine-grained annotations like word-level translation error labels (e.g. Fomicheva et al. 2021) are especially difficult to obtain. Obtaining automatic explanations to aid human annotators could boost their efficiency tremendously (e.g. Desmond et al. 2021). To our best knowledge, none of the current works has applied explainers in these settings for metrics. Semi-automatic labeling could also help with post-editing of translations (e.g. Specia and Farzindar 2010) to guide how sentences should be adapted. Goal 3 is important for metric developers that want to create such word-level translation error data sets and metric users that want to correct translations.
@@
Goal 4: Checking for social biases. Learned metrics might be biased towards certain types of texts used during training (e.g. Sun et al. 2022). Biases could be detected by observing explanations where sensitive attributes (for example, gender, political or racial aspects) are flagged to be important (Lipton 2016). For example, explanations could show that a metric considers male names more important for a translation than female names in the same scenario. Goal 4 can be important for all audiences. Affected users might be treated unfairly if biases occur in a metric. Metric users, metric developers, regulatory entities and managers might want to prevent this from happening to conform with given regulations.
@@ 3 Literature Review & Selection
To build the taxonomy in §4, we survey literature which either explicitly state that they address the explainability of MT metrics or propose techniques that make MT metrics or their outputs clearer to understand by certain audiences (techniques that provide explanations). We restrict our search and taxonomy to works that consider neural network based MT metrics. We conducted an initial literature search by selecting 11 “seed papers”. Then, we leveraged the API of Semantic Scholar999https://www.semanticscholar.org/ to query new papers citing the work of the seed papers and papers citing those papers (citation tree at depth 2). This amounted to 383 papers. Next, we filtered the papers by keywords, leaving 177 papers. Then, we manually annotated whether their contents fit our definitions. Thereby, we identified 12 papers that consider the explainability of MT metrics. In later stages, we progressively added further work released after the initial search and related works that we identified from other sources. In total, we describe 29 papers, some of which are not shown in Figure 4 due to space constraints.
Following §2.1, we consider higher granularity scores (e.g. word-level) as explanations for lower granularity metrics (e.g. sentence-level). However, without requiring faithfulness this definition allows to view any word-level metric as explainers for any sentence-level metric. Thereby, it would even include cases in which the word-level scores are determined completely independent from the sentence-level, i.e., cases where the explainer would not be faithful. Due to the high number of existing word-level approaches, for this explanation type, we only consider works that explicitly employ the word-level scores for explainability.
Likewise, following §2.1, metrics that report multiple quality aspects such as adequacy or fluency provide more explanatory value on why an explanation is good or bad than metrics that return a single score. For the quality aspect explanation type, we make an exemplary selection of techniques derived from recent metrics that leverage generation probabilities (e.g. Thompson and Post 2020; Yuan et al. 2021).

### v2

@@ 2.2 Explainability
Explanation type Further, we distinguish explanation types, i.e., the output format of the explanations (e.g., Wiegreffe and Marasovic, 2021). One example are natural language explanations that describe the model output or its internal decision process with free text. In the case of MT metrics, this could be a sentence like “The metric returned a score of 0.6 for the translation, as it correctly captures all details of the source sentence but has severe grammatical errors at positions […]”. There are many possible explanation types (e.g., Barredo Arrieta et al., 2020; Wiegreffe and Pinter, 2019), such as, feature importance explanations or counterfactuals. For our taxonomy, we select those that occur in our target papers (see §3). The definition of each explanation type is paired with the description of methods in §4.
@@
Model Access Most taxonomies of explainability techniques consider the required level of model access (e.g., Barredo Arrieta et al., 2020; Linardatos et al., 2021; Madsen et al., 2022). Here, techniques that require no access (that can work with any black box model) are called model-agnostic. Techniques that require access to model parameters are called model-specific, as only specific model architectures satisfy the requirements of the method. Due to the prevalent role of transformer architectures, many modern evaluation metrics use embeddings and attention mechanisms and allow for gradient computations. This allows to apply many techniques that are specific to these architectures to such metrics. A special type of model-specific explanations are those that are given by a model (here a metric) itself. We call these metrics “explainable by design”, as they are designed in a way that provides more details on their explanations or decision process. Examples for MT metrics in this category are unsupervised matching metrics like BERTScore Zhang et al. (2020a) and MoverScore Zhao et al. (2019), as they compute token-level similarity scores that can be used as explanations of the sentence-level score Leiter (2021); Fomicheva et al. (2021). Finally, there are also interpretable models which are easy to understand for humans such as Naive Bayes and decision trees Freitas (2014); Arya et al. (2019). We can normally obtain decision-understanding explanations of these models at the same time as predictions (for example, the corresponding path in the decision tree), while the models can also be considered explanations of themselves (for example, the trained decision tree itself). We do not use the dimension of interpretable models in our taxonomy (see §3) as none of the state-of-the-art metrics falls into this category. Model-agnostic techniques require least model access. Further, model-specific techniques, which use the gradient, embeddings or attention weights require more access. Finally, models that are explainable by design require access to the full model to generate explanations.666Note that besides interpretable models, none of the access levels inherently leads to faithful explanations, i.e., explanations that correctly describe the internal processes of a model. Hence, this property needs to be considered separately (e.g., Jacovi and Goldberg, 2020).
@@
Audiences Prior works in explainability point out that different audiences require different types of explanations (e.g., Barredo Arrieta et al., 2020; Sharma et al., 2020). For example, non-experts might have difficulties understanding explanations given in the domain language of experts. Barredo Arrieta et al. (2020) distinguish model users, affected users, developers, regulatory entities, and managers. Users of MT metrics are often experts in MT that want to evaluate or improve MT systems, i.e., MT developers. Another group of users are translation experts that want to compare MT systems and rate translations for usage in daily tasks (for example for post-editing). A third group of users might be non-experts that want to use the models to check their own translations or machine generated translations. Affected users might be the ones that receive translations that were graded by a metric or generated by a model that was supervised or tested with the metric. For example, Fernandes et al. (2022a) employ MT metrics in the decoding process of MT models. We consider users of those MT models as affected users. Developers are the designers of MT metrics. Regulatory entities and managers have an interest in enforcing guidelines. We describe the relation of the audiences towards goals of explainable MT metrics in §2.3 and associate audiences with explanation types in §4.
@@ 2.3 Goals of Explainable MT Metrics
In this section, we discuss goals that require specific focus for explainable MT evaluation. We orient ourselves on common goals proposed by Lipton (2016), Barredo Arrieta et al. (2020) and Jacovi and Goldberg (2021), and identify four specific ones that match the literature we identify in §3 and use cases that might commonly occur in our field of research (see Figure 2). Note that goal 1 and 4 address metric internals, while goal 2 and 3 mostly consider explanations of metric outputs (and with this, also the quality of the input translations). We follow with a short description of each of them below.
@@
Goal 1: Diagnose and improve metrics. By explaining why a metric predicted a certain score for a machine translation, where humans would assign a different score, developers might understand its weaknesses (e.g., Kaster et al., 2021; Sai et al., 2021). This can enable architectural changes or the selection of different data sets, leading to metric improvement. Likewise, explaining the whole metric can unveil whether it follows desired general properties or might otherwise be led astray by carefully crafted adversarial inputs. This goal requires explanations to be faithful to some extent, such that developers can understand the actual faults of the explained metric. Plausibility on the other hand may be harmful and trick developers into believing that their metric is working, therefore, it should be evaluated separately (Jacovi and Goldberg, 2020). Goal 1 is especially important for the audience metric developers, as they need to understand the shortcomings to improve on them.
@@
Goal 2: Make metrics more expressive and accessible. A metric that assigns a score on a sentence-level is difficult to understand, especially for users who are not experienced with MT and MT metrics. If further information about a metric’s output is provided, such as which words it considers incorrect, the metric may become more accessible and thus foster trust (which may lead to more widespread adoption of the metric). Exemplary works in this direction are Fomicheva et al. (2021) and Bao et al. (2023). Accessible metrics might also help with use cases like translation error correction in the downstream tasks of language learning (non-experts) or human translation services (translation experts) (e.g., Zouhar et al., 2023). In these latter two use cases, a high average explanation plausibility (to translation experts) and a strong metric performance are important. This is because, in cases where (1) the metric output is correct and (2) the explanation of the output is plausible to an expert, the explanation very likely correctly describes the shortcomings and strengths of the translation (on the level of an expert).777To foster trust, it might be sufficient, but unethical, if explanations of any metric (also weak ones) are plausible to its users (not necessarily to experts) (e.g., Jacovi and Goldberg, 2020; Jin et al., 2023). The problem is that users could be tricked to believe into the performance of weak metrics or adapt the mindset of wrong explanations that have believable arguments. Hence, benchmarks of metrics and explanations in their correlation to human expert quality scores and explanations are important. Generally, users should be informed that hallucinations, where an explanation is untrustworthy, might occur. Faithfulness is not necessary in this scenario, because the quality of the explanation in describing the translation already follows from (1) and (2). However, faithfulness is still desirable, as the internal decision process of a strong metric might support giving correct explanations (see §2.2). Goal 2 is important for the audience of metric users. These could, for example, be non-experts that want to understand specific metric outputs or developers of MT systems that want to understand if a metric is a good fit for their tasks (see Figure 1).
@@
Goal 3: Support Semi-Automatic labeling. Fine-grained (human) annotations like word-level translation error labels (e.g., Fomicheva et al., 2021) are especially difficult to obtain. Obtaining automatic explanations to aid human annotators could boost their efficiency tremendously (e.g., Desmond et al., 2021). To our best knowledge, none of the current works has applied explainers in these settings for metrics. Semi-automatic labeling could also help with post-editing of translations (e.g., Specia and Farzindar, 2010) to guide how sentences should be adapted. Goal 3 is important for metric developers that want to create such word-level translation error data sets and metric users who want to correct translations. The requirements for faithfulness and plausibility are the same as for goal 2.
@@
Goal 4: Checking for social biases. Learned metrics might be biased towards certain types of texts used during training (e.g., Sun et al., 2022). Biases could be detected by observing explanations where sensitive attributes (for example, gender, political or racial aspects) are flagged to be important (Lipton, 2016). For example, explanations could show that a metric considers male names more important for a translation than female names in the same scenario. Goal 4 can be important for all audiences. Affected users might be treated unfairly if biases occur in a metric. Metric users, metric developers, regulatory entities and managers might want to prevent this from happening to conform with given regulations. The requirements for faithfulness and plausibility are the same as for goal 1.
@@ 3 Literature Review & Selection
To build the taxonomy in §4, we survey literature which either explicitly state that they address the explainability of MT metrics or propose techniques that make MT metrics or their outputs clearer to understand by certain audiences (techniques that provide explanations). We restrict our search and taxonomy to works that consider neural network based MT metrics. We conducted an initial literature search by selecting 11 “seed papers”. Then, we leveraged the API of Semantic Scholar888https://www.semanticscholar.org/ to query new papers citing the work of the seed papers and papers citing those papers (citation tree at depth 2). This amounted to 383 papers. Next, we filtered the papers by keywords, leaving 177 papers. Then, we manually annotated whether their contents fit our definitions. Thereby, we identified 12 papers that consider the explainability of MT metrics. In later stages, we progressively added further work released after the initial search and related works that we identified from other sources. In total, we describe 37 papers, some of which are not shown in Figure 3 due to space constraints, but discussed throughout the chapter.
Following §2.1, we consider higher granularity scores (e.g., word-level) as explanations for lower granularity metrics (e.g., sentence-level), i.e., feature importance explanations. Due to the high number of existing word-level approaches, for this explanation type we only consider works that explicitly employ word-level scores to explain sentence-level scores.
Likewise, following §2.1, metrics that report multiple quality aspects such as adequacy or fluency provide more explanatory value on why a translation is good or bad than metrics that return a single score. For the quality aspect explanation type, we make an exemplary selection of techniques derived from recent metrics that leverage generation probabilities (e.g., Thompson and Post, 2020; Yuan et al., 2021).